Sophie received a $32,750 loan from a bank that was charging interest at 4.50% compounded semi-annually.

a. How much does she need to pay at the end of every 6 months to settle the loan in 4 years?

Round to the nearest cent

b. What was the amount of interest charged on the loan over the 4-year period?

Round to the nearest cent
Answer #1

Know tempound interest formula is, d nt A = Pr (its where pe A future amount Poinita cipal amount date of interest n - No. ofbs Now, Here, A = $ 39130.72 and p = $32750 So, Interest charged is given by I = A-P I 391 30.72 32750 =) 1 = 6380. 72 charge